# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ 3.2 Main result: $\mathtt{RMSProp}$ ($\beta_{1}=0,\beta_{2}\in[0,1]$)
We can only upper bound $\left\|\frac{{\bm{g}}_{t}}{\sqrt{{\bm{v}}_{t}+\epsilon}}\right\|_{\infty}^{2}$ by $\frac{1}{1-\beta_{2}}$ without more fine-grained analysis on the relationship between gradients in each step, which will greatly hurt the dependence of convergence rate on $1-\beta_{2}$. However, even though the update at step $t$ for one specific coordinate $i$ can be as large as $\frac{1}{\sqrt{1-\beta_{2}}}$ with some very large $g_{t,i}$, the average moving speed for each coordinate should be close to $1$. Therefore, we introduce a slightly stronger definition in Definition 3.3, which allows us to decompose the second order term into each coordinate according to Lemma 3.12. It also facilitates the analysis for the coordinate-wise first order term. We note this definition also appears in Assumption 2.3 of the concurrent work Maladkar et al. (2024).
@@
For any $\mathbf{H}=(H_{1},\ldots,H_{d})\in\mathbb{R}^{d}$, we say a function $L$ is $\mathbf{H}$-smooth coordinate-wisely w.r.t. $\ell_{\infty}$ norm , iff for any $i\in[d]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$, $\left|\nabla_{i}L({\bm{x}})-\nabla_{i}L({\bm{y}})\right|\leq H_{i}\left\|{\bm{x}}-{\bm{y}}\right\|_{\infty}.$
By definition, $\mathbf{H}$-smoothness coordinate-wisely w.r.t. $\ell_{\infty}$ norm implies $\sum_{i=1}^{d}H_{i}$ smoothness w.r.t. $\ell_{\infty}$ norm. We also need 3.4 to measure the influence of noise in stochastic setting.
@@
There exist constants $\sigma_{i}$ such that
$\mathbb{E}\left(\nabla_{i}L_{t}({\bm{x}})-\nabla_{i}L({\bm{x}})\right)^{2}\leq\sigma_{i}^{2}$
for any $i\in[d]$, $t\in\mathbb{N}$ and ${\bm{x}}\in\mathbb{R}^{d}$.
Due to the limitation of space, we only present the main result here. The sketch of the proof is presented in Section 3.4. We present the complete proof for the generalized blockwise Adam algorithm in Appendix C. The proof incorporates some key steps from Li & Lin (2024), extending them to accommodate the generalized algorithm and different smoothness assumptions.
@@
$F=2\ln\left(1+\frac{\sum_{i=1}^{d}\sigma_{i}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\infty}^{2}+\max_{i\in[d]}H_{i}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
We can determine the convergence rate of RMSprop by applying appropriate hyperparameters on Theorem 3.5. The optimal hyperparameters $\eta$ and $\beta_{2}$ can be selected by minimizing $E$. We would assume that $v_{0}+\epsilon>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ and $\frac{1}{1-\beta_{2}}=\mathtt{poly}(T)$. Then we can simplify the term by considering $F=O(\log T)$.
The two terms involving $\sum_{i=1}^{d}\sigma_{i}$ have a lower bound $\Theta\left(\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{2}}\right)$, which can reached by $1-\beta_{2}=\Theta\left(\frac{\log{T}}{T}\right)$. With this choice of $1-\beta_{2}$, the three terms involving $\eta$ has a lower bound $\Theta\left(\sqrt{\frac{\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}}{T}}\right)$ reached by $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$. Such choices of hyperparameters can give the optimal convergence rate for stochastic case in Corollary 3.6. For convenience, we define $R\triangleq\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}$, which will be the core term in Corollaries 3.6 and 3.7.
@@
Let $\{L_{t}\}_{t=1}^{T}$ be independent stochastic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Theta(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}+\delta_{T}\right).$
with $\delta_{T}=\sqrt{\frac{dv_{0}}{T(1-\beta_{2})}}\exp{\left(-\frac{T(1-\beta_{2})}{8}\right)}\left[\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}\right]$
Here $\delta_{T}$ can be smaller than any polynomial of $T$ by manipulating the value of $\frac{T(1-\beta_{2})}{\log{T}}=\Theta(1)$. Then $\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}$ is the leading term w.r.t. $T$ in the rate. The coefficient only involves $\sum_{i=1}^{d}\sigma_{i}$, suggesting that the rate can be much improved when noise is small. Below we get the convergence rate with the same hyperparameters in deterministic case in Corollary 3.7.
@@
Let $\{L_{t}\}_{t=1}^{T}$ be deterministic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Omega(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\max_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ for any polynomial $\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\delta_{T}\right).$
@@
Corollary 3.7 almost recovers Theorem 3.2, except the smoothness constant here $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{(1,1)}$ is worse than that in Theorem 3.2, which is $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{\infty}$, because it always holds that $\|\cdot\|_{1,1}\geq\|\cdot\|_{\infty}$. This gap is due to a technical difficulty of analyzing $\mathtt{Adam}$ or $\mathtt{RMSProp}$, as mentioned in the beginning of Section 3.2.
@@
While many previous works rely on the relatively large magnitude of $\epsilon$ compared to ${\bm{v}}_{t}$ and give a bound in the regime of $\mathtt{SGD}$ when the adaptive effect is dominated by the constant $\epsilon$ (Zaheer et al., 2018; De et al., 2018), our result actually prefers $\epsilon$ to be 00 while maintaining the value of $v_{0}+\epsilon$. We also note the dependence of our bound in Theorem 3.5 on $v_{0}$ is very mild and logarithmic. Theorem 3.5 has similar convergence rate for all $v_{0}$ of magnitude at most $\mathtt{poly}(T)$, while most previous result only addresses the case where $v_{0,i}$ is at the scale of noise (Li & Lin, 2024) or 00. The main reason for this adaptivity to a wide range of $v_{0}$ is our specific choice of $\beta_{2}=1-\Theta(\frac{\log T}{T})$, which allows the initial large $v_{0}$ to decay fast and resume normal training. Other existing results using $\beta_{2}=1-\Theta(1/T)$ (Défossez et al., 2022; Li & Lin, 2024) cannot allow large initial value $v_{0}$ because $v_{0}$ only decays a constant fraction throughout the training and the effective learning rate will be too small.
@@ 3.3 A unified analysis for blockwise $\mathtt{Adam}$
In this subsection, we present convergence analysis for a broader class of adaptive algorithms defined in Algorithm 3, which could be thought as a coarser version of $\mathtt{Adam}$. It does pre-conditioning blockwisely (specified by a partition function $\Phi:[d]\rightarrow[B]$ where $B$ is the number of blocks) instead of coordinate-wisely. Since $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ can be viewed as special cases blockwise $\mathtt{Adam}$ (Algorithm 3) with $\Phi_{\mathtt{Adam}}:i\mapsto i$ and $\Phi_{\mathtt{AdaSGD}}:i\mapsto 1$ respectively, any convergence results for Algorithm 3 would imply convergence of $\mathtt{Adam}$ and $\mathtt{AdaSGD}$. Finally we also note that such blockwise $\mathtt{Adam}$ has been recently studied empirically by some concurrent work, where the algorithm is named by $\mathtt{Adamini}$ (Zhang et al., 2024b) and $\mathtt{Adalayer}$ (Zhao et al., 2024).
@@
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
@@
   $v_{t,b}\leftarrow\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\sum_{B(i)=b}g_{t,i}^{2}\right)/d_{b}$
   $x_{t,i}\leftarrow x_{t-1,i}-\eta_{t}\frac{m_{t,i}}{\sqrt{v_{t,B(i)}+\epsilon}}$
@@
We first introduce some more notations. We define the vector ${\bm{x}}_{(b)}$ as $[x_{i}]_{\Phi(i)=b}$ and the submatrix $\mathbf{A}_{(b),(b^{\prime})}$ as $[A_{i,j}]_{\Phi(i)=b,\Phi(j)=b^{\prime}}$.
@@
We define the $(\infty,2)$-norm w.r.t. partition $\Phi$ of vector ${\bm{x}}$ as the $\ell_{\infty}$ norm of the vector $\left(\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}\right)_{b=1}^{B}$, which is $\max_{b\in[B]}\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}$. For convenience, we will denote it by $\left\|{\bm{x}}\right\|_{\Phi}$ or just call it $\Phi$-norm. We denote its dual norm by $\|{\bm{x}}\|_{\Phi,*}$, which is equal to $\sum_{b=1}^{B}\sqrt{d_{b}}\left\|{\bm{x}}_{(b)}\right\|_{2}$.
Definition 3.9 (Generalized version of Definition 3.3).
For any partition function $\Phi:[d]\rightarrow[B]$ and $\mathbf{H}=(H_{1},\ldots,H_{B})\in\mathbb{R}^{B}$, we say a function $L$ is $\mathbf{H}$-smooth blockwisely w.r.t. $(\infty,2)$ norm , iff for any $b\in[B]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$,
$\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}})-\nabla_{(b)}L({\bm{y}})\right\|_{2}\leq H_{b}\left\|{\bm{x}}-{\bm{y}}\right\|_{\Phi}$
We further define the $\Phi$-smoothness of loss $L$ by $H(L,\Phi)=\sum_{b=1}^{B}H_{b}$, where $\{H_{b}\}_{b=1}^{B}$ are the smallest numbers making $L$ is $\mathbf{H}$-smooth blockwisely in the above sense.
We note that the above defined blockwise $(\infty,2)$-smoothness is both a generalization of the coordinate-wise smoothness defined in Definition 3.1 (corresponding to the case of each block only containing $1$ coordinate) and the standard $\ell_{2}$ smoothness (corresponding to the case of only having one block). In the former case, we have $B=d$, $\Phi_{\mathtt{Adam}}$ is the identity mapping $i\mapsto i$ and it holds that $H(L,i\mapsto i)\geq\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}\geq\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{\infty}$. In the latter case, we have $B=1$, $\Phi_{\mathtt{AdaSGD}}$ is the mapping $i\mapsto 1$ and $H(L,i\mapsto 1)=\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$.
Similar to the coordinate-wise case, one can show that $\sum_{b=1}^{B}H_{b}$ w.r.t. partition $\Phi$ is an upper bound for the smoothness of loss $L$ w.r.t. to $\Phi$-norm.

### v3

@@ 3.2 Main result: $\mathtt{RMSProp}$ ($\mathtt{Adam}$ with $\beta_{1}=0,\beta_{2}\in[0,1]$)
We can only upper bound $\left\|\frac{{\bm{g}}_{t}}{\sqrt{{\bm{v}}_{t}+\epsilon}}\right\|_{\infty}^{2}$ by $\frac{1}{1-\beta_{2}}$ without more fine-grained analysis on the relationship between gradients in each step, which will greatly hurt the dependence of convergence rate on $1-\beta_{2}$. However, even though the update $\frac{g_{t,i}}{\sqrt{v_{t,i}+\epsilon}}$ at step $t$ can be as large as $\frac{1}{\sqrt{1-\beta_{2}}}$ with some very large $g_{t,i}$, the average moving speed for each coordinate should be close to $1$. Therefore, we introduce Definition 3.3, which is slightly stronger than Definition 3.1 but allows us to decompose the second order term into each coordinate according to Lemma 3.13. It also facilitates the coordinate-wise analysis for the first order term. We note this definition also appears in Assumption 2.3 of the concurrent work Maladkar et al. (2024).
@@
For any $\mathbf{H}=(H_{1},\ldots,H_{d})\in\mathbb{R}^{d}$, we say a function $L$ is $\mathbf{H}$-smooth coordinate-wisely w.r.t. $\ell_{\infty}$ norm , if and only if $\left|\nabla_{i}L({\bm{x}})-\nabla_{i}L({\bm{y}})\right|\leq H_{i}\left\|{\bm{x}}-{\bm{y}}\right\|_{\infty}$ for any $i\in[d]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$.
(1,1)-norm as a surrogate complexity measure for coordinate-wise smoothness.
$H_{i}$ in Definition 3.3 is determined by $\sup_{{\bm{x}}}\sum_{j=1}^{d}\left|\nabla^{2}_{i,j}L({\bm{x}})\right|$, which is difficult to compute because it requires taking supreme over the entire domain. A computationally-tractable alternative is to approximate $\sum_{i=1}^{d}H_{i}$ locally by the $(1,1)$-norm of Hessian of loss along the training trajectory. We are the first to provide an efficient approximation algorithm with concentration guarantees in Section D.2, which uses hessian-vector product against random Cauchy vectors.
By definition, $\mathbf{H}$-smoothness coordinate-wisely w.r.t. $\ell_{\infty}$ norm implies $\sum_{i=1}^{d}H_{i}$-smoothness w.r.t. $\ell_{\infty}$ norm. We also need 3.4 to measure the influence of noise in the stochastic setting.
@@
There exist constants $\sigma_{i}$ such that $\mathbb{E}\left[\nabla_{i}L_{t}({\bm{x}})-\nabla_{i}L({\bm{x}})\right]^{2}\leq\sigma_{i}^{2}$ for any $i\in[d]$, $t\in\mathbb{N}$ and ${\bm{x}}\in\mathbb{R}^{d}$.
We present the main result in Theorem 3.5. The sketch of the proof is presented in Section 3.4 and the complete proof for the generalized blockwise Adam algorithm is presented in Appendix C. The proof incorporates some key steps from Li & Lin (2024), extending them to accommodate the generalized algorithm and different smoothness assumptions.
@@
$F=2\ln\left(1+\frac{\sum_{i=1}^{d}\sigma_{i}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\infty}^{2}+\sum_{i\in[d]}H_{i}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
We can determine the convergence rate of RMSprop by choosing appropriate hyperparameters $\eta$ and $\beta_{2}$ in Theorem 3.5 to minimize $E$. We would assume that $v_{0}+\epsilon>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\sum_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ and $\frac{1}{1-\beta_{2}}=\mathtt{poly}(T)$. Then we can simplify the term by considering $F=O(\log T)$.
The two terms involving $\sum_{i=1}^{d}\sigma_{i}$ have a lower bound $\Theta\left(\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{2}}\right)$, which is achieved with $1-\beta_{2}=\Theta\left(\frac{\log{T}}{T}\right)$. Then the three terms involving $\eta$ has a lower bound $\Theta\left(\sqrt{\frac{\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}}{T}}\right)$ reached by $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$. These hyperparameter choices yield the optimal convergence rate for stochastic case in Corollary 3.6. For convenience, we define $R\triangleq\left(L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})\right)\sum_{i=1}^{d}H_{i}$, which will be the core term in Corollaries 3.6 and 3.7.
@@
Let $\{L_{t}\}_{t=1}^{T}$ be independent stochastic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Theta(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\sum_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}+\delta_{T}\right)$
with $\delta_{T}=\sqrt{\frac{dv_{0}}{T(1-\beta_{2})}}\exp{\left(-\frac{T(1-\beta_{2})}{8}\right)}\left[\left(\frac{R}{T}\right)^{\frac{1}{4}}+\sqrt{\sum_{i=1}^{d}\sigma_{i}}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}\right]$.
Here $\delta_{T}$ can be smaller than any polynomial of $T$ by manipulating the value of $\frac{T(1-\beta_{2})}{\log{T}}=\Theta(1)$. Then $\sum_{i=1}^{d}\sigma_{i}\left(\frac{\log{T}}{T}\right)^{\frac{1}{4}}$ is the leading term w.r.t. $T$ in the rate whose coefficient only involves $\sum_{i=1}^{d}\sigma_{i}$. It suggests that the rate can be much improved when noise is small. Below we get the convergence rate with the same hyperparameters in deterministic case in Corollary 3.7.
@@
Let $\{L_{t}\}_{t=1}^{T}$ be deterministic losses satisfying 3.4 and that their expectation $L$ is ${\bm{H}}$-coordinate-wisely smooth w.r.t. $\ell_{\infty}$ norm. For $\beta_{1}=0$, $1-\beta_{2}=\Omega(\frac{\log{T}}{T})$, $\epsilon=0$, $\eta=\Theta\left(\sqrt{\frac{L({\bm{x}}_{0})-\min_{\bm{x}}L({\bm{x}})}{T\sum_{i=1}^{d}H_{i}}}\right)$ and $v_{0}>(\sum_{i=1}^{d}\sigma_{i}^{2}+\|\nabla L({\bm{x}}_{0})\|_{\infty}^{2}+\sum_{i}H_{i}^{2}\eta^{2})/\mathtt{poly}(T)$ for any polynomial $\mathtt{poly}(T)$, we have that
$\min_{\frac{T}{2}<t\leq T}\left\|{\bm{g}}_{t}\right\|_{1}=O\left(\sqrt{\frac{R}{T}}+\delta_{T}\right)$
@@
Corollary 3.7 almost recovers Theorem 3.2, except for the smoothness constant. Specifically, it uses $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{1,1}$ which is larger than $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{\infty}$ in Theorem 3.2 as $\|\cdot\|_{1,1}\geq\|\cdot\|_{\infty}$ always holds. This gap is due to technical difficulty of analyzing $\mathtt{Adam}$ as mentioned in the beginning of Section 3.2.
@@
While many previous works rely on the relatively large magnitude of $\epsilon$ compared to ${\bm{v}}_{t}$ and give a bound in the regime of $\mathtt{SGD}$ when the adaptive effect is dominated by the constant $\epsilon$ (Zaheer et al., 2018; De et al., 2018), our result actually prefers $\epsilon$ to be 00 while maintaining the value of $v_{0}+\epsilon$. We also note the dependence of our bound in Theorem 3.5 on $v_{0}$ is very mild and logarithmic. Theorem 3.5 has similar convergence rates for all $v_{0}$ of magnitude at most $\mathtt{poly}(T)$, while most previous result only addresses the case where $v_{0,i}$ is at the scale of noise (Li & Lin, 2024) or 00. The main reason for this adaptivity to a wide range of $v_{0}$ is our specific choice of $\beta_{2}=1-\Theta(\frac{\log T}{T})$, which allows the initial large $v_{0}$ to decay fast and resume normal training. Other existing results using $\beta_{2}=1-\Theta(1/T)$ (Défossez et al., 2022; Li & Lin, 2024) cannot allow large initial value $v_{0}$ because $v_{0}$ only decays a constant fraction throughout the training and the effective learning rate will be too small.
@@ 3.3 A unified analysis for blockwise $\mathtt{Adam}$
In this subsection, we present convergence analysis for a broader class of adaptive algorithms defined in Algorithm 3. It can be viewed as a coarser version of $\mathtt{Adam}$ because it does pre-conditioning blockwisely instead of coordinate-wisely. Since $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ can be viewed as special cases of blockwise $\mathtt{Adam}$ (Algorithm 3) with $\Phi_{\mathtt{Adam}}:i\mapsto i$ and $\Phi_{\mathtt{AdaSGD}}:i\mapsto 1$ respectively, any convergence results for Algorithm 3 would imply convergence of $\mathtt{Adam}$ and $\mathtt{AdaSGD}$. Finally we also note that such blockwise $\mathtt{Adam}$ has been recently studied empirically by some concurrent work, where the algorithm is named by Adam-mini (Zhang et al., 2024b) and Adalayer (Zhao et al., 2024).
@@
initialization ${\bm{x}}_{0}$, stochastic losses $\{L_{t}\}_{t=1}^{T}$
@@
   $v_{t,b}\leftarrow\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\sum_{\Phi(i)=b}g_{t,i}^{2}\right)/d_{b}$
   $x_{t,i}\leftarrow x_{t-1,i}-\eta_{t}\frac{m_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}$
@@
We first introduce more notations. For a partition function $\Phi:[d]\rightarrow[B]$ where $B$ is the number of blocks, $(b)$ is defined as $\Phi^{-1}(b)=\{i|\Phi(i)=b\}$ and $d_{b}=\#(b)$, the number of parameters in block $b$. We define the vector ${\bm{x}}_{(b)}$ as $[x_{i}]_{\Phi(i)=b}$ and the submatrix $\mathbf{A}_{(b),(b^{\prime})}$ as $[A_{i,j}]_{\Phi(i)=b,\Phi(j)=b^{\prime}}$.
@@
We define the $(\infty,2)$-norm w.r.t. partition $\Phi$ of the vector ${\bm{x}}$ as the $\ell_{\infty}$ norm of the vector $\left(\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}\right)_{b=1}^{B}$, which is $\max_{b\in[B]}\frac{\left\|{\bm{x}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}}$. For convenience, we will denote it by $\left\|{\bm{x}}\right\|_{\Phi}$ or just call it $\Phi$-norm. We denote its dual norm by $\|{\bm{x}}\|_{\Phi,*}$, which is equal to $\sum_{b=1}^{B}\sqrt{d_{b}}\left\|{\bm{x}}_{(b)}\right\|_{2}$.
We will need Definition 3.9 to characterize the smoothness of the loss function when analyzing general blockwise $\mathtt{Adam}$. We note $\Phi_{\mathtt{Adam}}$-smoothness also appears in Assumption 2 in Bernstein et al. (2018) to analyze $\mathtt{SignGD}$. Previous and concurrent work (Ene et al., 2021; Liu et al., 2024; Jiang et al., 2024) employ the same assumption to analyze AdaGrad. We are the first to analyze $\mathtt{Adam}$ under such smoothness assumption. Moreover, we are the first to empirically measure it for loss functions in real task with theoretical guarantee.
Definition 3.9 ($\Phi$-smoothness).
We say a diagonal matrix ${\bm{A}}$ follows partition $\Phi$ if and only if the diagonal elements of ${\bm{A}}$ are constant within each block defined by $\Phi$, i.e., there are $a_{1},\cdots,a_{B}\in\mathbb{R}$, such that ${\bm{A}}_{i,i}=a_{\Phi(i)}$ for any $i\in[d]$.
We say a twice-differentiable $L$ is $H$-smooth under partition $\Phi$ if there exists a diagonal matrix ${\bm{A}}$ following partition $\Phi$ such that $H=\operatorname{Tr}({\bm{A}})$ and ${\bm{A}}$ dominates $\left|\nabla^{2}L({\bm{x}})\right|$ for all $x$. We further define the $\Phi$-smoothness of loss $L$, denoted by $H(L,\Phi)$, as the smallest constant $H$ such that $L$ is $H$-smooth under parition $\Phi$, that is,
$\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}H(L,\Phi)=\min_{\begin{subarray}{c}{\bm{A}}\text{ follows }\Phi\\ {\bm{A}}\succeq\left|\nabla^{2}L({\bm{x}})\right|,\forall{\bm{x}}\end{subarray}}\operatorname{Tr}({\bm{A}})$ (1)
We note that $\Phi$-smoothness is different from the smoothness under $\Phi$-norm, where the latter is equal to $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}\left\|\nabla^{2}L({\bm{x}}){\bm{u}}\right\|_{\Phi,*}$. For each ${\bm{x}}$, it holds that
$\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}\left\|\nabla^{2}L({\bm{x}}){\bm{u}}\right\|_{\Phi,*}=\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}\sup_{\left\|{\bm{v}}\right\|_{\Phi}\leq 1}{\bm{v}}^{\top}\nabla^{2}L({\bm{x}}){\bm{u}}=\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}\left|{\bm{u}}^{\top}\nabla^{2}L({\bm{x}}){\bm{u}}\right|.$
When ${\bm{A}}\succeq\left|\nabla^{2}L({\bm{x}})\right|$, we have that $\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}\left|{\bm{u}}^{\top}\nabla^{2}L({\bm{x}}){\bm{u}}\right|\leq\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}{\bm{u}}^{\top}{\bm{A}}{\bm{u}}$. When diagonal ${\bm{A}}$ follows partition $\Phi$, we have that $\sup_{\left\|{\bm{u}}\right\|_{\Phi}\leq 1}{\bm{u}}^{\top}{\bm{A}}{\bm{u}}=\operatorname{Tr}({\bm{A}})$. So $\Phi$-smoothness defined in Definition 3.9 is always no smaller than the smoothness under $\Phi$-norm.
Another advantage with the definition is that the $H(L,\Phi)$ will always non-increase with a more fine-grained partition. For two partition functions $\Phi_{1}$ and $\Phi_{2}$, we say $\Phi_{1}$ includes $\Phi_{2}$ if and only if $\Phi_{1}(i)=\Phi_{1}(j)$ for any $i,j\in[d]$ such that $\Phi_{2}(i)=\Phi_{2}(j)$. If a diagonal matrix ${\bm{A}}$ follows partition $\Phi_{1}$ and partition $\Phi_{1}$ includes $\Phi_{2}$, then ${\bm{A}}$ also follows $\Phi_{2}$. So we have that $\{{\bm{A}}\mid{\bm{A}}\text{ follows }\Phi_{1},{\bm{A}}\succeq\left|\nabla^{2}L({\bm{x}})\right|\text{ for all }{\bm{x}}\}\subseteq\{{\bm{A}}\mid{\bm{A}}\text{ follows }\Phi_{2},{\bm{A}}\succeq\left|\nabla^{2}L({\bm{x}})\right|\text{ for all }{\bm{x}}\}$ and $H(L,\Phi_{1})$ is no smaller than $H(L,\Phi_{2})$. Since $\Phi_{\mathtt{AdaSGD}}$ includes any partition and any partition includes $\Phi_{\mathtt{Adam}}$, $H(L,\Phi_{\mathtt{AdaSGD}})$ is the largest and $H(L,\Phi_{\mathtt{Adam}})$ is the smallest among all the partitions $\Phi$.
With generalized 3.10 on noise, we can prove the result for blockwise $\mathtt{Adam}$ in Theorem 3.11.